OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

fog(x)=f(g(x)) =f(8-x) =(8-x)^2-2(8-x)+9 =(64+x2-16x)-16+2x+9 =x2-14x+57 now fog(-4)= put the value of x as -4...

OpenStudy (anonymous):

Please ask in the math section.

OpenStudy (trancenova):

Hi please post questions in the correct section :)